# Bosch China Investment Ltd

### v4

@@ 4.4 Ablation Study
To prove the effectiveness of our purposed components. We first evaluate the performance of OPME,  OPMD and GMIF. Later, go into the details of the OPME and OPMD. As shown in Tab. 3, we evaluate the efficacy of our order-preserving mechanism in OPME and OPMD, as well as the global interaction design in GMIF, by progressively scaling down our model. We do this by gradually removing OPME, OPMD, and GMIF to establish a baseline model, and then conducting message hiding training using COCO with a 4 BPP payload. In the absence of these components 2D convolution layers are utilized. It becomes apparent that the absence of both OPME and OPMD, as well as GMIF, hinders the model’s ability to effectively hide the secret message within the cover image. A model employing only GMIF can enhance accuracy through more efficient message-image fusion. However, the inclusion of OPME and OPMD substantially improves message recovery accuracy, underscoring the importance of preserving the sequential order of the message during encoding.
As illustrated in Tab. 4, we evaluate the effectiveness of internal modules of OPME and OPMD at 4 BPP using the COCO dataset. We begin by testing models containing only an MLP and gradually add MHSA and PEs. The inclusion of solely PEs or MHSA after the MLP module yields suboptimal performances. The optimal performance is achieved with the default design of our OPME and OPMD, which incorporates MLP, MHSA, and PEs.
To further validate whether OPME and OPMD effectively model the independent nature of message segments, we analyze the ratio between the cumulative sum of the top n eigenvalues and the sum of all eigenvalues of PEs [Wang and Chen(2020)]. Specifically, we compare this characteristic of PEs obtained from ViT with those from OPME. As shown in Fig. 6, positional embeddings from ViT display significant eigenvalues, suggesting shared positional information among numerous image patches for image classification tasks. In contrast, PEs from OPME lack prominent eigenvalues, indicating much greater independence of positional information. This demonstrates that OPME effectively assigns distinct positional information to individual message segments.
MLP MHSA PE ACC PSNR SSIM
$\checkmark$ $\times$ $\times$ 98.59 40.59 0.9826
$\checkmark$ $\checkmark$ $\times$ 98.68 40.70 0.9833
$\checkmark$ $\times$ $\checkmark$ 97.92 40.05 0.9786
$\checkmark$ $\checkmark$ $\checkmark$ 99.24 41.78 0.9875
Table 4: Ablation of components of OPME and OPMD.
@@ 5 Conclusion
We propose StegaFormer, a novel message hiding framework for high accuracy message concealment and recovery. StegaFormer is able to increase the message recovery accuracy by emphasizing the order of message bits and cross-modality feature fusion. Experiments on real-world datasets demonstrate that our method consistently outperforms the SOTA methods in terms of message recovery accuracy, message capacity, and imperceptibility.
Supplementary Material for Effective Message Hiding with Order Preserving Mechanisms
This appendix is organized as follows.
A Kernel Size Experiment
Figure 7: The effects of different kernel size to the recovery accuracy and imperceptibility.
As depicted in Fig. 7, we conduct several experiments related to the kernel size of the CNN-based approach to investigate its limitations. The experiments for SteganoGAN [Zhang et al.(2019a)Zhang, Cuesta-Infante, and Veeramachaneni] is based on the officially released code. We choose the base model of SteganoGAN as it avoids the complexities of other variants that involve embedding messages at multiple convolutional stages, making it difficult to discern how message features are encoded into the image features. Since there is no official code released for ChatGAN [Tan et al.(2021)Tan, Liao, Liu, Cao, and Jiang], our experiments for this approach is based on our own implementations. We integrate the channel attention module into the SteganoGAN as described in the paper of ChatGAN. All models are trained using the DIV2K dataset with 4 BPP message payload for 20 epochs, with all other configurations remaining unchanged, except for the kernel size of these two models.
B Selection of Query, Key and Value for GMIF
[…]
Query Key and Value ACC PSNR SSIM
MSG MSG+IM 50.00% 168.8 1.0
MSG IM 99.23% 41.35 0.9863
IM MSG+IM 99.27% 41.87 0.9877
IM MSG 99.22% 41.43 0.9863
MSG+IM MSG+IM 99.07% 41.40 0.9862
Table 5: The comparison between different configurations of GMIF. MSG and IM represent the message feature and image feature respectively. MSG+IM denotes the summation of image and message features. The underline indicates the configuration adopted in our paper.
C Detailed Configurations of StegaFormer
The message concealment pipeline comprises three OPME modules for order-preserving message encoding. Each OPME consists of one MHSA, with 2 heads. Additionally, there are three GMIF modules to fuse the message-image features. Each GMIF includes one W-MSA, with the number of heads set to 2 and the window size set to 16. Regarding the message recovery pipeline, there are four W-MSA modules, all configured to 16 window size and 2 attention heads. Furthermore, one OPMD serving as the message head for recovering the secret message in the message recovery process. The number of heads is set to 2 for MHSA inside OPMD. The detailed configuration of the message recovery pipeline is listed in Tab. 6.
The number of parameters in StegaFormer is solely determined by the length of message segment $L_{\text{ms}}$. As mentioned in the main paper, the typical settings for $L_{\text{ms}}$ are 16, 32, 48, and 64. On the other hand, increasing the range of message element $N_{\text{r}}$ does not affect the number of parameters in StegaFormer. The typical number of parameters for our StegaFormer at different values of $L_{\text{ms}}$ are listed in Tab. 7. We also list the Flops and number of parameters at typical 1 BPP in Tab. 8.
There are two approaches to increase the message hiding capacity in our method: increasing the length of message segments $L_{\text{ms}}$ and expanding the range of message elements $N_{r}$. When $N_{r}=1$, typical choices for $L_{\text{ms}}$ include 16, 32, 48, and 64, corresponding to message capacities of 1, 2, 3, and 4 BPP, respectively. $N_{r}$ can also be increased to achieve higher message capacities. For example, $N_{r}=2^{i}-1,\quad\text{where }i=2,3,4$, represents 2, 3, and 4 bits per message element, respectively. Combining these two factors can lead to a higher message capacity for our models.
We list the experiment results related to different configurations of $L_{\text{ms}}$ and $N_{\text{r}}$ we have tested in searching for maximum message capacity in Tab. 9 and underlined the listed combinations in the main paper. All the experiments are trained with COCO dataset.
Module Name Sub Module Number of Layers Number of Heads Window Size Output Feature Dimension
W-MSA 1 NA 1 2 16 $2\times L_{\text{ms}}$
W-MSA 2 NA 1 2 16 $4\times L_{\text{ms}}$
W-MSA 3 NA 1 2 16 $8\times L_{\text{ms}}$
W-MSA 4 NA 1 2 16 $8\times L_{\text{ms}}$
OPMD MHSA 1 2 NA $8\times L_{\text{ms}}$
MLP 1 NA NA $L_{\text{ms}}$
Table 6: Detailed configurations for message recovery pipeline. The number added to W-MSA represents the order of W-MSA in the message recovery process from right to left. NA denotes not available.
$(L_{\text{ms}},N_{r})$ Capacity Encoder Decoder $(L_{\text{ms}},N_{r})$ Capacity Encoder Decoder $(L_{\text{ms}},N_{r})$ Capacity Encoder Decoder
(16, 1) 1 BPP 10.75M 2.39M (48, 1) 3 BPP 48.25M 10.70M (48, 3) 6 BPP 48.25M 10.70M
(32, 1) 2 BPP 21.46M 4.76M (64, 1) 4 BPP 85.75M 19.00M (32, 15) 8 BPP 21.46M 4.76M
Table 7: Number of parameters of StegaFormer from 1 to 8 BPP message capacity. M denotes a million parameters.
Method FLOPs (G) # Param (M)
StegaFormer 74.3 13.14
Message Concealment of StegaFormer 58.7 10.75
Message Recovery of StegaFormer 15.6 2.39
Table 8: Flops and number of parameters of StegaFormer in 1 BPP message capacity.
Configuration list Configuration list
($L_{\text{ms}}$,$N_{\text{r}}$) BPP ACC PSNR SSIM ($L_{\text{ms}}$,$N_{\text{r}}$) BPP ACC PSNR SSIM
(16,1) 1 99.95% 47.83 0.9969 (16,7) 3 98.72% 42.86 0.9890
(32,1) 2 99.85% 45.30 0.9943 (32,7) 6 95.03% 36.72 0.9673
(48,1) 3 99.68% 43.37 0.9914 (48,7) 9 78.45% 35.50 0.9432
(64,1) 4 99.27% 41.87 0.9877 (16,15) 4 96.32% 40.33 0.9799
(16,3) 2 99.61% 44.98 0.9934 (32,15) 8 91.78% 34.70 0.9508
(32,3) 4 97.46% 41.37 0.9847 (48,15) 12 66.31% 30.37 0.8673
(48,3) 6 95.65% 40.37 0.9803
Table 9: The experiment results of StegaFormer using different configurations of $L_{\text{ms}}$ and $N_{\text{r}}$. The selected configurations are underlined.
D Detailed Configurations of Experiments
The training process of our model consists of 100,000 iterations, with a batch size of 2. We set $\lambda_{1}=1\times 10^{-4}$ and $\lambda_{2}=1\times 10^{-6}$ to balance different losses. For comparison, we set the image size to $256\times 256$ for all models. All other settings remain as in the officially released models. We utilize the official models of SteganoGAN and LISO [Chen et al.(2023)Chen, Kishore, and Weinberger] since they are publicly available. We use our own inplamentation of ChatGAN based on SteganoGAN since there is no officially released code.
E More Qualitative Examples
More qualitative comparisons of residual and stego image generated by StegaFormer and other two methods are in Fig. 8. Additionally, more comparisons of residual and stego images generated by StegaFormer at different message capacities are in Fig. 9.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$I_{cover}$ (a) $I_{stego}$ (b) $I_{stego}$ (c) $I_{stego}$ (a) $I_{res}$ (b) $I_{res}$ (c) $I_{res}$
Figure 8: Qualitative results of 3 BPP message hiding using different methods: (a) SteganoGAN, (b) ChatGAN and (c) Ours. $I_{cover}$, $I_{stego}$ and $I_{res}$ represent cover image, stego image and residual image respectively. To enhance visualization, the residual images are multiplied by 5. Best viewed in digital version.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) (b) (c) (d) (e) (f) (g)
Figure 9: Qualitative results of StegaFormer: (a) cover images $I_{cover}$, (b) 1 BPP, (c) 2 BPP, (d) 3 BPP, (e) 4 BPP, (f) 6 BPP, and (g) 8 BPP. Rows 1, 3, 5, and 7, along with columns (b) to (g), represent the stego images $I_{stego}$. Rows 2, 4, 6, and 8, along with columns (b) to (g), represent the residual images $I_{res}$. To enhance visualization, the residual images are multiplied by 5. Best viewed in digital version.
F Experiments of Steganalysis
As described in Sec. 4.3. We conduct steganalysis experiments using SiaStegNet [You et al.(2020)You, Zhang, and Zhao]. Among all the experiments, we use COCO dataset for this experiments. We generate 25,000 training sets consisting of cover-stego image pairs and 500 testing pairs using listed approaches and our model. SiaStegNet is trained for 2 epochs. We do not continue training SiaStegNet for more epochs since the Siamese approach is highly effective when cover and stego image pairs are available for training. Training for more than 2 epochs would result in a 100% detection rate for all the methods. The aim of this experiment is to demonstrate the superiority of security of our approach.

### v5

@@ 4.4 Ablation Study
To prove the effectiveness of our purposed components. We first evaluate the performance of OPME, OPMD and GMIF. Later, go into the details of the OPME and OPMD.
4.4.1 Effectiveness of OPME , OPMD and GMIF
As illustrated in  Tab.˜3, to assess the effectiveness of our order-preserving in OPME and OPMD and global interaction design in GMIF, we scale down our model by removing the OPME and OPMD and GMIF step by step to set up a baseline model, then perform the message hiding training at a 4 BPP message payload. In their absence, these components are replaced with 2D convolution layers. It is evident that the absence of both OPME and OPMD and GMIF prevents the model from successfully hiding the secret message within the cover image. A model using only GMIF can boost accuracy via more effective message-image fusion. The involvement of OPME and OPMD significantly increases the accuracy of message recovery, highlighting the effectiveness of preserving message sequential order in message encoding.
Models ACC PSNR SSIM
baseline 50.01% 53.4 0.9988
+GMIF 87.60% 43.78 0.9908
+GMIF+OPME+OPMD 99.24% 41.78 0.9875
Table 3: Ablation of the proposed OPME, OPMD and GMIF. We replace our proposed components with 2D convolution layers in the baseline model which disable the global message-image fusion and the order-preserving message encoding and decoding, respectively.
4.4.2 Effectiveness of PE within OPME and OPMD
MLP MSA PE ACC PSNR SSIM
$\checkmark$ $\times$ $\times$ 98.59% 40.59 0.9826
$\checkmark$ $\checkmark$ $\times$ 98.68% 40.70 0.9833
$\checkmark$ $\times$ $\checkmark$ 97.92% 40.05 0.9786
$\checkmark$ $\checkmark$ $\checkmark$ 99.24% 41.78 0.9875
Table 4: Ablation of the proposed OPME and OPMD. We compare the performance of different version of them, which gradually add the PE and MSA to validate the effectiveness of these components.
As illustrated in  Tab.˜4, we assess the efficacy of the features of OPME and OPMD at 4 BPP using COCO dataset. We start with the model only contains MLP as a message encoder and decoder. The addition of solely positional embedding or MSA after the MLP module results in sub-optimal performance. The best performance is achieved with the default design of our OPME and OPMD, which include MLP, MSA, and PEs.
To further validate whether OPME and OPMD appropriately models the independent nature of message segments, we assess the positional embeddings learned by OPME. More specifically, we compare the characteristics of positional embeddings obtained from ViT with those from OPME. As illustrated in  Fig.˜7, the positional embeddings obtained from ViT exhibit a prevalence of significant eigenvalues, suggesting that numerous image patches share similar positional information for image classification task. In contrast, the positional embeddings acquired from OPME do not exhibit prominent eigenvalues, indicating that the positional embeddings in OPME contain more independent positional information than those in ViT. This demonstrates that OPME has effectively learned to assign distinct positional information to each individual message segment.
@@ 5 Conclusion
In this paper, we propose StegaFormer, a novel message hiding framework for high accuracy message concealment and recovery. StegaFormer is able to significantly increase the message recovery accuracy by emphasizing the order of message bits and cross-modality feature fusion. Experiments on real-world datasets demonstrate that our method consistently outperforms the state-of-the-art methods in terms of message recovery accuracy, message capacity, and imperceptibility.
Ethical Statement
There are no ethical issues.
Acknowledgments
We thank Wang Wenfeng for conducting early experiments and Lu Lvjian for reviewing the manuscript.